A random variable B is distributed as P{B< x} = 1 − e^-(ax) (x > 0).
Find density function of a) sqrt(B) , b) (1/a) ln{B}
1
Expert's answer
2012-11-16T05:12:26-0500
An empty 210 kg elevator accelerates upward
at 1.8 m/s2.
The acceleration of gravity is 9.8 m/s2 .
What is the tension in the cable that lifts
the elevator cab?
Answer in units of N

According to the second Newton's law the tension in the cable that lifts the elevator cab is

T = M*g + M*a = M*(g+a) = 210[kg]*(9.8[m/s&sup2;] + 1.8[m/s&sup2;]) = 2436[N].